# Spiral renormalization group flow and universal entanglement spectrum of the non-Hermitian 5-state Potts model

### v1

@@ Abstract
The $1+1$D quantum $Q$-state Potts model is a paradigmatic example of this phenomenon. This model is known to undergo a continuous phase transition for $Q\leq 4$ and a first-order transition for $Q>4$, consistent with the two real fixed points for $Q\leq 4$, a critical and tricritical one, which overlap at $Q=4$. The weakly first-order nature of the $Q=5$ case has been conjectured to originate from a complex conformal field theory (CCFT) at the complex fixed points that governs the walking regime [10, 9]. This scenario has been theoretically proposed based on the analytic continuation of Coulomb gas techniques. Recently, a lattice regularization of the 5-state Potts CCFT was found by extending the Potts Hamiltonian with a non-Hermitian term, verifying the existence of the complex fixed points $\mathcal{C}$ and $\overline{\mathcal{C}}$ [30]. The deformed 5-state Potts model is given by
@@
where $\hat{X}$ is the spin shift operator which acts on the $Q$ states labeled |$\ket{n},n=0,\ldots,Q-1$ as $\hat{X}\ket{n}=\ket{(n+1)\mod Q}$ and $\hat{Z}$ the phase operator which acts as $\hat{Z}\ket{n}=e^{\frac{2\pi ni}{Q}}\ket{n}$. The critical value $\lambda_{c}\in\mathbb{C}$ was found by exact diagonalization methods to simulate the complex fixed points [30]. The RG flow around these fixed points, as predicted in [9] and verified in Fig. 1, materializes as a logarithmic spiral in the complex theory space, along with a straight line representing the walking behavior of the original 5-state Potts model.
Figure 1: The flow of the running coupling parameter $g_{\varepsilon^{\prime}}$ acquired through the cost function fitting [30, 26] (a) for $\mathcal{\overline{C}}$ and (b) for $\mathcal{C}$ around their respective fixed points for system sizes $L=16-24$. The perturbation parameter at $\lambda_{c}$ is denoted by a blue dot, while red arrows show the walking behavior of values on the real axis.
@@
Tensor networks—We define the ground state wavefunction of the deformed Potts Hamiltonian as being the eigenstate corresponding to the eigenvalue with the smallest real part, and assume that the Schmidt coefficients decay fast enough such that it can be represented as a matrix product state (MPS) [33]. In particular, we impose the $\mathbb{Z}_{5}$ subgroup symmetry of the Hamiltonian for computational efficiency and to identify the charges of the CCFT primary operators easily. We use an MPS ansatz with open BC (and a Hamiltonian with periodic BC, hence with one long-ranged interaction term) such that can employ the standard density matrix renormalization group (DMRG) [35]. The excited states are determined by an algorithm that targets the quasiparticle excitations [32, 12]. We find ground states with a bond dimension $\chi=400$, while for excited states we use $\chi=600$, both with an algorithmic tolerance of $10^{-6}$. We simulate up to system sizes $L=28$, assuring that we are only in the finite-size scaling regime and that second-order terms remain negligible [22]. Due to the small non-Hermitian deformation, the singular values of the MPS retain their ordering, explaining why the MPS representations obtained through DMRG and the Quasiparticle ansatz remain robust [26]. Only in the imaginary results, a known problem due to the high bond dimension needed for the real part [27, 26], does it make slight systematic errors. We use the Julia packages TensorKit.jl and MPSKit.jl to perform the numerical simulations [11, 31].
@@
Conformal data—To increase the precision of the fixed point $\lambda_{c}$, we consider leading irrelevant perturbations from the operators $\varepsilon^{\prime}$, $T\overline{T}$, $T^{2}$ and $\overline{T}^{2}$ allowed by the symmetries of the Hamiltonian. The (spatial) symmetry requirements leave only $\partial^{4}\mathbb{I},\overline{\partial}^{4}\mathbb{I}$ and $\Box\varepsilon^{\prime}$ which have vanishing one-point functions or orthogonal Verma modules to our group of operators [5]. We simulate the Hamiltonian with PBC to exploit radial quantization. Consequently, at the critical point where $g_{\varepsilon^{\prime}}=0$, the ground state energy and energy gaps depend on the central charge $c$ and scaling dimensions $\Delta_{n}$ of the CCFT according to [1, 4]
[…]
@@
where $x=1/L$, $v$ is the speed of light, $f$ is the ground state energy density and $a,\tilde{a}_{n}$ take into account the leading irrelevant perturbations. For each point $\lambda$ in theory space, we first recover the non-universal parameter $v$ by plugging in the theoretical central charge $c$ and ground state energies for $L=8-14$. We extrapolate the fit of Eq. (S.5) for all pairs of subsequent data points to $L\to+\infty$ to obtain $v\approx 2.880-0.705i$. These intermediate system sizes allow us to exclude sub-subleading finite-size corrections, while also avoiding deviations due to the finite bond dimension. Now we evaluate the deviation between the leading order contribution to $\Delta_{n}$ in Eq. (S.5) and the theoretical value $\Delta_{n}^{\text{th}}$ [9]:
@@
We now extract the conformal data with the refined critical parameter. To do so, we fit Eq. (S.5) to the fit parameters $\Delta_{n}$, $f$, $c$, $a$ and $\tilde{a}_{n}$ for $L=8-14$. The conformal scaling dimension estimates are visible in Table 1. We compare these values with the theoretical predictions found in [9]. Additionally, we include the scaling dimensions extracted through exact diagonalization (ED) as done in [30] for $\lambda_{c}^{\text{prev}}=0.079+0.060i$ and $L=8-13$. We observe that the improved accuracy of our new fixed point yields more accurate estimates for the conformal dimensions. The ED result at $\lambda_{c}^{\text{prev}}$ only has better accuracy compared to the MPS algorithms at $\lambda_{c}$ for the imaginary parts of the conformal data. This is most likely due to the non-Hermitian MPS ansatz having more untrustworthy low entanglement contributions [26]. Moreover, the tensor network framework deteriorates with increasing scaling dimension due to deviations from the area law. Lastly, the imaginary part of $\Delta_{\varepsilon}$ in Table 1 was fitted for $1/L^{2}$, which according to Fig. 2 does not appear to be correct. The scaling which is closest to the theoretical value is $1/L^{1.05}$ with a value of $\mathrm{Im}(\Delta_{\varepsilon})=-0.2245(1)i$. It remains unclear whether this observation has a deeper meaning or if it is due to the same simulation artifacts discussed above.
[…]
$c$ $1.1375-0.0211i$ $1.1376(1)-0.0212(1)i$ $1.1405-0.0224i$
$\sigma$ $0.1336-0.0205i$ $0.1336(1)-0.0205(1)i$ $0.1334-0.0206i$
$\sigma^{\prime}$ $1.1107-0.1701i$ $1.1105(1)-0.1697(1)i$ $1.1081-0.1702i$
$\partial\sigma$ $1.1336-0.0205i$ $1.1334(1)-0.0203(1)i$ $1.1328-0.0206i$
$\varepsilon$ $0.4656-0.2245i$ $0.4656(1)-0.2240(1)i$ $0.4633-0.2240i$
$\partial\varepsilon$ $1.4656-0.2245i$ $1.4650(1)-0.2207(5)i$ $1.4623-0.2218i$
$\varepsilon^{\prime}$ $1.9083-0.5987i$ $1.9081(2)-0.5964(2)i$ $1.8998-0.5977i$
Table 1: The central charge and conformal dimensions of the lowest-lying operators of the $Q=5$ Potts CCFT. From left to right: the theoretical predictions [9], the values from DMRG or QPA fitted to Eq. (S.5) for $\lambda_{c}=0.0788+0.0603i$ for system sizes $L=8-14$, and the values determined by [30] at $\lambda_{c}^{\text{prev}}=0.079+0.060i$ and system sizes $L=8-13$ using exact diagonalization.
@@
Figure 3: The entanglement spectrum of the ground state in the (a) zero sector and (b) a charged sector of $\mathbb{Z}_{5}$ for the deformed 5-state Potts Hamiltonian at $\lambda_{c}=0.0788+0.0603i$ for system sizes $L=12-28$ and periodic boundary conditions. The dashed lines and red numbers are the theoretical energy levels and their degeneracies, respectively [29]. We fit the data to their theoretical values for the first two energy levels of the zero sector.
@@
$\rho_{RL}=\frac{\ket{\psi_{L}}\bra{\psi_{R}}}{\braket{\psi_{R}}{\psi_{L}}},$ (6)
where we identify |$\ket{\psi_{L}}=\ket{\psi_{R}}^{*}$ due to the transpose symmetry of the Hamiltonian. In particular, we analyze the entanglement Hamiltonian $K_{\mathcal{C}}$ [6] of the CCFT, which we propose remains related to the density matrix as
@@
For large enough system sizes, the entanglement spectrum of the lattice with PBC is expected to flow to the free-free conformal boundary condition (CBC), since it is the only one that respects the $S_{Q}$ symmetry of the Hamiltonian. The boundary CCFT operator content for the free-free CBC has been found in [29] through finite-size scaling. In particular, three conformal towers with primaries $\mathbb{I}$, $\phi_{3,1}$, and $\phi_{5,1}$ were confirmed with degeneracies one, four, and eleven, respectively [24]. For our simulations, where we impose the $\mathbb{Z}_{5}$ symmetry on the MPS, we expect the identity and $\phi_{3,1}$ operator to be located in the zero sector and four charged sectors of $\mathbb{Z}_{5}$, respectively. The final low-lying operator can be uniquely divided into a three-fold degenerate operator in the zero sector and a two-fold degenerate one in the four charged sectors, as dictated by its possible $S_{5}$ restriction to $\mathbb{Z}_{5}$. This eleven-dimensional operator most likely couples to both a five- and the six-dimensional irreducible representation (sector) of $S_{5}$ as those indeed yield the observed restriction to the sectors of $\mathbb{Z}_{5}$.
@@
B B.   Fixed Point Precision
@@
[…]
@@
Figure S1: The drift of the critical value due to subleading finite-size corrections on the $g_{\varepsilon^{\prime}}$ minimum of the cost function, Eq. (S.6), for system sizes $L=10-24$. We note that the minimum at $L=18$ and $L=19$ overlap. We perform a fit of $y=ax^{2}+b$ with $x=1/L$ for all data points and for $L=17-24$. Their zero points $b$ are marked by a cross and yield an estimate for the critical point, denoted $\lambda_{CF}$. The estimate $\lambda_{c}=0.0788+0.0603i$ is also included.
@@
To combat the subleading FSE, we extrapolate the drift to $L\to+\infty$ and determine a more accurate critical point. Figure S1 shows this process. By using our data across system sizes $L=10-24$, we find a value of $\lambda_{CF}=0.0786+0.0608i$. However, we find that irrelevant terms with larger scaling dimensions persist for smaller system sizes up to $L\sim 13$. In larger systems, these terms disappear, and the fitting thereof $L=17-24$ yields a different estimate $\lambda_{CF}=0.0788+0.0605i$.
@@
We do not overfit solely by considering the largest system sizes, because in this case, the drift exactly matches the size of the discrete grid used in theoretical space. In Figure S1 we also included our newly proposed critical point $\lambda_{c}$, which was determined purely based on finite-size scaling theory with subleading corrections. We see that this is very close to our last estimate of the cost function, with even $\mathrm{Re}(\lambda_{c})=Re(\lambda_{CF})$. This suggests that the system sizes under consideration are large enough to exclude subleading FSE. The mismatch in the imaginary part remains unclear, but could be attributed to either sub-subleading FSE or the MPS ansatz making less accurate approximations in the imaginary axis, as shown in Section E.
@@
We present the resulting RG flow in Figure S4. This figure illustrates, to some extent, the expected topology of the RG flow as presented in [1]. The values that fitted the real parameters on the real axis were $x=0.076-0.015i$ and $y=0.245-0.057i$. This is, to our knowledge, the first RG flow construction of the deformed 5-state Potts model in a (1+1)-dimensional quantum system. Another CCFT RG flow has previously been identified [3]. However, this RG flow has a different topology as it is based on a (2+0)D statistical analysis of the O$(n>2)$ model, which is in turn related to the $Q>4$ Potts model. This relation causes its shape to change, not making it comparable to the theoretical predictions of [1]. Regardless, there are areas where the flow seems to deviate in Figure S4. For example, at large $\lambda$ values, it is not even possible to keep real parameters on the real axis. This is most likely due to the failure of the simulation method, as discussed in Section E. Additionally, due to the limited search grid of $\lambda$ used, we did not capture the entire RG flow.
Figure S4: The evolution of $\lambda_{w}=x+y\cdot g_{\varepsilon^{\prime}}$ for $L=16-24$ where $x$ and $y$ are fixed such that real theories reside and remain on the real axis. The points corresponding to $\lambda_{c}=0.0788\pm 0.0603$ and the evolution of values on the real axis are in blue and red respectively.
D D.   Open Boundary Conditions
@@
Figure S5: The entanglement spectrum of the ground state in the (a) zero sector and (b) a charged sector of $\mathbb{Z}_{5}$ for the deformed 5-state Potts Hamiltonian at $\lambda_{c}=0.0788+0.0603i$ for system sizes $L=16-64$ and open boundary conditions. Here we use DMRG with $\chi=600$. The dashed lines and red numbers are the theoretical energy levels and their degeneracies, respectively [7]. We fit only the first two energy levels of the zero sector to their theoretical values.
E E.   Sources of error
@@
Working within the MPS manifold leads to different sources of errors. First, the MPS representation has a finite bond dimension. This introduces deviations from the true wavefunction because only finite entanglement is present. As a result, there is a perturbation from the true critical point [Huang2024]. This is particularly problematic for critical systems, where the ground state’s correlation length diverges, and for excited states, which may violate the area law of entanglement. The limited entanglement at the MPS boundary also introduces additional boundary effects. However, we confirmed that DMRG and QPA reach their bond saturation within the algorithmic tolerance for $\chi\geq 400$ and $600$, respectively, for system sizes up to $L=24$, regardless of the additional difficulties associated with simulating non-Hermitian models. For these systems, another limitation occurs. The ordering of the singular values used in the DMRG algorithm is no longer a direct representation of the non-Hermitian entanglement from $\rho_{RL}$. The true entanglement spectrum consists of the eigenvalues $e_{i}$ of the density matrix, which we calculate for an MPS using the method explained in the main text. To investigate how well this MPS ansatz can find the true entanglement structure of non-Hermitian systems, we look at the difference between the singular values $s_{i}$ of the ground state wavefunction |$\ket{\psi_{R}}$ and eigenvalues of the corresponding density matrix. In Hermitian systems, these are related by $s_{i}^{2}=e_{i}$. In non-Hermitian systems, the singular values will remain real, and the eigenvalues can become complex. We show in Figure S6 (a) the square of the singular values of the ground state wavefunction |$\ket{\psi_{R}}$, together with the singular and eigenvalues of the density matrix $\rho_{RL}$, of the ground state at the middle of an $L=24$ chain at the critical point $\lambda_{c}=0.0788+0.0603i$. These are ordered by largest real part. We also plot the absolute value of the real and imaginary parts of every eigenvalue. Furthermore, this graph is nearly identical throughout the bulk, indicating that the boundary effects have little influence on the entanglement structure.
Firstly, it is visible that the squares of the singular values of |$\ket{\psi_{R}}$ and $\rho_{RL}$ are nearly identical. This is because the two are the same up to a normalization constant. More interestingly, we show that the density matrix is not guaranteed to be positive semi-definite as it possesses negative real eigenvalues. We observe this by the increase in absolute value around the value of 250, regardless of its largest real ordering. The largest negative real part of an eigenvalue ends up being of order $10^{-6}$. It is also interesting to note that the ordering according to the largest real eigenvalues corresponds (mostly) to the ordering of the largest imaginary part, indicating some correspondence between the real and imaginary parts. Due to the observed disconnect between singular and eigenvalues, it remains unknown how much entanglement is actually being neglected by the MPS approximation. This is partly rectified by the equivalent ordering property, assuring that if we discard singular values, it will be of similar order in both real and imaginary parts. Since we aim for a certain tolerance in the real spectrum, this requires a certain bond dimension $\chi$. However, as we increase $\chi$, it provides more imaginary contributions to the singular values, which are, relative to their real counterpart, smaller in absolute value. As a result, by imposing a large bond dimension, a larger portion of low, untrustworthy, and imaginary entanglement is retained. This might explain why the imaginary entanglement is known to increase incorrectly with the bond dimension [6], and why the conformal dimensions and spectra shown in the main text and in Figure S5 are less accurate in the imaginary part.
Figure S6: The singular value decomposition $(s_{i})$ and (absolute magnitude) of the eigenvalues $(e_{i})$ of the non-Hermitian density matrix $\rho_{RL}$ and ground state |$\ket{\psi_{R}}$, ordered by (a) largest real part and (b) largest magnitude, of the deformed 5-state Potts model at $\lambda_{c}$ in the middle of an $L=24$ chain.
@@
To recapitulate, in this work we employed DMRG and the Quasiparticle ansatz to determine the ground state and excited states, respectively [9, 2]. We found these to agree with ED up to the imposed algorithmic tolerance of $10^{-6}$ for $\lambda_{c}=0.0788+0.0603i$ and $L=10$. Only when increasing the imaginary part of $\lambda$ in the deformed Potts Hamiltonian do we observe that these algorithms make errors that are multiple orders of magnitude larger than the imposed tolerance. This might be due to the smallest real ordering, used in both algorithms, no longer representing the ground state. Moreover, these algorithms do not account for the differences between left and right eigenvectors. We briefly mention that we also attempted to find excited states via DMRG by projecting out the ground state. We found these to agree with QPA for $\chi\geq 400$, but convergence was much slower beyond the first excited state.

### v2

@@ Abstract
The 1+11+1D quantum $Q$-state Potts model is a paradigmatic example of this phenomenon. This model is known to undergo a continuous phase transition for $Q\leq 4$ and a first-order transition for $Q>4$, consistent with the two real fixed points for $Q\leq 4$, a critical and tricritical one, which overlap at $Q=4$. The weakly first-order nature of the $Q=5$ case has been conjectured to originate from a complex conformal field theory (CCFT) at the complex fixed points that governs the walking regime [11, 10]. This scenario has been theoretically proposed based on the analytic continuation of Coulomb gas techniques. Recently, a lattice regularization of the 5-state Potts CCFT was found by extending the Potts Hamiltonian with a non-Hermitian term, verifying the existence of the complex fixed points $\mathcal{C}$ and $\overline{\mathcal{C}}$ [30]. The deformed 5-state Potts model is given by
@@
[…]
Figure 1: The flow of the running coupling parameter $g_{\varepsilon^{\prime}}$ acquired through the cost function fitting [30, 26] (a) for $\mathcal{\overline{C}}$ and (b) for $\mathcal{C}$ around their respective fixed points for system sizes L=1624L=16-24. The perturbation parameter at $\lambda_{c}$ is denoted by a blue dot, while red arrows show the walking behavior of values on the real axis.
@@
Tensor networks—We define the ground state wavefunction of the deformed Potts Hamiltonian as being the eigenstate corresponding to the eigenvalue with the smallest real part, and assume that the Schmidt coefficients decay fast enough such that it can be represented as a matrix product state (MPS) [33]. In particular, we impose the $\mathbb{Z}_{5}$ subgroup symmetry of the Hamiltonian for computational efficiency and to identify the charges of the CCFT primary operators easily. We use an MPS ansatz with open BC (and a Hamiltonian with periodic BC, hence with one long-ranged interaction term) such that can employ the standard density matrix renormalization group (DMRG) [35]. The excited states are determined by an algorithm that targets the quasiparticle excitations [32, 12]. We find ground states with a bond dimension $\chi=400$, while for excited states we use $\chi=600$, both with an algorithmic tolerance of $10^{-6}$. We simulate up to system sizes $L=64$, assuring that we are only in the finite-size scaling regime and that second-order terms remain negligible [22]. Due to the small non-Hermitian deformation, the singular values of the MPS retain their ordering, explaining why the MPS representations obtained through DMRG and the Quasiparticle Ansatz (QPA) remain robust. As further analyzed in [26], the algorithms only experience slight systematic errors in the imaginary results, a known problem [27]. We use the Julia packages TensorKit.jl and MPSKit.jl to perform the numerical simulations [9, 31].
@@
Conformal data—To increase the precision of the fixed point $\lambda_{c}$, we consider leading irrelevant perturbations from the operators $\varepsilon^{\prime}$, $T\overline{T}$, $T^{2}$ and $\overline{T}^{2}$ allowed by the symmetries of the Hamiltonian The (spatial) symmetry requirements leave only $\partial^{4}\mathbb{I},\overline{\partial}^{4}\mathbb{I}$ and $\Box\varepsilon^{\prime}$ which have vanishing one-point functions or orthogonal Verma modules to our group of operators [5]. We simulate the Hamiltonian with PBC to exploit radial quantization. Consequently, at the critical point where $g_{\varepsilon^{\prime}}=0$, the ground state energy and energy gaps depend on the central charge $c$ and scaling dimensions $\Delta_{n}$ of the CCFT according to [1, 4]
$\begin{aligned} \frac{E_{0}}{L}&=f-\frac{\pi vc}{6}x^{2}+ax^{4},\\ (E_{n}-E_{0})L&=2\pi v\Delta_{n}+\tilde{a}_{n}x^{2},\end{aligned}$ (3)
@@
where $x=1/L$, $v$ is the speed of light, $f$ is the ground state energy density and $a,\tilde{a}_{n}$ take into account the leading irrelevant perturbations. For each point $\lambda$ in theory space, we first recover the non-universal parameter $v$ by plugging in the theoretical central charge $c$ and ground state energies for L=814L=8-14. We extrapolate the fit of Eq. (S.5) for all pairs of subsequent data points to $L\to+\infty$ to obtain $v\approx 2.880-0.705i$. These intermediate system sizes allow us to exclude sub-subleading finite-size corrections, while also avoiding deviations due to the finite bond dimension. Now we evaluate the deviation between the leading order contribution to $\Delta_{n}$ in Eq. (S.5) and the theoretical value $\Delta_{n}^{\text{th}}$ [10]:
@@
We now extract the conformal data with the refined critical parameter. To do so, we fit Eq. (S.5) to the fit parameters $\Delta_{n}$, $f$, $c$, $a$ and $\tilde{a}_{n}$ for L=814L=8-14. The conformal scaling dimension estimates are visible in Table 1. We compare these values with the theoretical predictions found in [10]. Additionally, we include the scaling dimensions extracted through exact diagonalization (ED) as done in [30] for $\lambda_{c}^{\text{prev}}=0.079+0.060i$ and L=813L=8-13. We observe that the improved accuracy of our new fixed point yields more accurate estimates for the conformal dimensions. The ED result at $\lambda_{c}^{\text{prev}}$ only has better accuracy compared to the MPS algorithms at $\lambda_{c}$ for the imaginary parts of the conformal data. This is most likely due to the non-Hermitian MPS ansatz having more untrustworthy low entanglement contributions [26]. Moreover, the tensor network framework deteriorates with increasing scaling dimension due to deviations from the area law. Lastly, the imaginary part of $\Delta_{\varepsilon}$ in Table 1 was fitted for $1/L^{2}$, which according to Fig. 2 does not appear to be correct. The scaling which is closest to the theoretical value is $1/L^{1.05}$ with a value of $\mathrm{Im}(\Delta_{\varepsilon})=-0.2245(1)i$. It remains unclear whether this observation has a deeper meaning or if it is due to the same simulation artifacts discussed above.
Complex CFT $\text{DMRG/QPA at }\lambda_{c}$ ED at $\lambda_{c}^{\text{prev}}$
$c$ 1.13750.0211i1.1375-0.0211i 1.1376(1)0.0212(1)i1.1376(1)-0.0212(1)i 1.14050.0224i1.1405-0.0224i
$\sigma$ 0.13360.0205i0.1336-0.0205i 0.1336(1)0.0205(1)i0.1336(1)-0.0205(1)i 0.13340.0206i0.1334-0.0206i
$\sigma^{\prime}$ 1.11070.1701i1.1107-0.1701i 1.1105(1)0.1697(1)i1.1105(1)-0.1697(1)i 1.10810.1702i1.1081-0.1702i
$\partial\sigma$ 1.13360.0205i1.1336-0.0205i 1.1334(1)0.0203(1)i1.1334(1)-0.0203(1)i 1.13280.0206i1.1328-0.0206i
$\varepsilon$ 0.46560.2245i0.4656-0.2245i 0.4656(1)0.2240(1)i0.4656(1)-0.2240(1)i 0.46330.2240i0.4633-0.2240i
$\partial\varepsilon$ 1.46560.2245i1.4656-0.2245i 1.4650(1)0.2207(5)i1.4650(1)-0.2207(5)i 1.46230.2218i1.4623-0.2218i
$\varepsilon^{\prime}$ 1.90830.5987i1.9083-0.5987i 1.9081(2)0.5964(2)i1.9081(2)-0.5964(2)i 1.89980.5977i1.8998-0.5977i
Table 1: The central charge and conformal dimensions of the lowest-lying operators of the $Q=5$ Potts CCFT. From left to right: the theoretical predictions [10], the values from DMRG or QPA fitted to Eq. (S.5) for $\lambda_{c}=0.0788+0.0603i$ for system sizes L=814L=8-14, and the values determined by [30] at $\lambda_{c}^{\text{prev}}=0.079+0.060i$ and system sizes L=813L=8-13 using exact diagonalization.
@@
Figure 3: The entanglement spectrum of the ground state in the (a) zero sector and (b) a charged sector of $\mathbb{Z}_{5}$ for the deformed 5-state Potts Hamiltonian at $\lambda_{c}=0.0788+0.0603i$ for system sizes L=1264L=12-64 and periodic boundary conditions. The dashed lines and red numbers are the theoretical energy levels and their degeneracies, respectively [29]. We fit the data to their theoretical values for the first two energy levels of the zero sector.
@@
[…]
@@
For large enough system sizes, the entanglement spectrum of the lattice with PBC is expected to flow to the free-free conformal boundary condition (CBC), since it is the only one that respects the $S_{Q}$ symmetry of the Hamiltonian. The boundary CCFT operator content for the free-free CBC has been found in [29] through finite-size scaling. In particular, three conformal towers with primaries $\mathbb{I}$, $\phi_{3,1}$, and $\phi_{5,1}$ were confirmed with degeneracies one, four, and eleven, respectively [24]. For our simulations, where we impose the $\mathbb{Z}_{5}$ symmetry on the MPS, we expect the identity and $\phi_{3,1}$ operator to be located in the zero sector and four charged sectors of $\mathbb{Z}_{5}$, respectively. The final low-lying operator can be uniquely divided into a three-fold degenerate operator in the zero sector and a two-fold degenerate one in the four charged sectors, as dictated by its possible $S_{5}$ restriction to $\mathbb{Z}_{5}$ This eleven-dimensional operator most likely couples to both a five- and the six-dimensional irreducible representation (sector) of $S_{5}$ as those indeed yield the observed restriction to the sectors of $\mathbb{Z}_{5}$.
@@
G B. Fixed Point Precision
@@
$\begin{aligned} \frac{E_{0}}{L}&=f-\frac{\pi vc}{6}x^{2},\\ (E_{n}-E_{0})\times L&=2\pi v\Delta_{n},\end{aligned}$ (S.5)
@@
Figure S1: The drift of the critical value due to subleading finite-size corrections on the $g_{\varepsilon^{\prime}}$ minimum of the cost function, Eq. (S.6), for system sizes L=1024L=10-24. We note that the minimum at $L=18$ and $L=19$ overlap. We perform a fit of $y=ax^{2}+b$ with $x=1/L$ for all data points and for L=1724L=17-24. Their zero points $b$ are marked by a cross and yield an estimate for the critical point, denoted $\lambda_{CF}$. The estimate $\lambda_{c}=0.0788+0.0603i$ is also included.
@@
To combat the subleading FSE, we extrapolate the drift to $L\to+\infty$ and determine a more accurate critical point. Figure S1 shows this process. By using our data across system sizes L=1024L=10-24, we find a value of $\lambda_{CF}=0.0786+0.0608i$. However, we find that irrelevant terms with larger scaling dimensions persist for smaller system sizes up to $L\sim 13$. In larger systems, these terms disappear, and the fitting thereof L=1724L=17-24 yields a different estimate $\lambda_{CF}=0.0788+0.0605i$.
@@
We do not overfit solely by considering the largest system sizes, because in this case, the drift exactly matches the size of the discrete grid used in theoretical space. In Figure S1 we also included our newly proposed critical point $\lambda_{c}$, which was determined purely based on finite-size scaling theory with subleading corrections. We see that this is very close to our last estimate of the cost function, with even $\mathrm{Re}(\lambda_{c})=Re(\lambda_{CF})$. This suggests that the system sizes under consideration are large enough to exclude subleading FSE. The mismatch in the imaginary part remains unclear, but could be attributed to either sub-subleading FSE or the MPS ansatz making less accurate approximations in the imaginary axis, as shown in Section D D.
@@
We present the resulting RG flow in Figure S4. This figure illustrates, to some extent, the expected topology of the RG flow as presented in [1]. The values that fitted the real parameters on the real axis were x=0.0760.015ix=0.076-0.015i and y=0.2450.057iy=0.245-0.057i. This is, to our knowledge, the first RG flow construction of the deformed 5-state Potts model in a (1+1)-dimensional quantum system. Another CCFT RG flow has previously been identified [3]. However, this RG flow has a different topology as it is based on a (2+0)D statistical analysis of the O(n>2)(n>2) model, which is in turn related to the $Q>4$ Potts model. This relation causes its shape to change, not making it comparable to the theoretical predictions of [1]. Regardless, there are areas where the flow seems to deviate in Figure S4. For example, at large $\lambda$ values, it is not even possible to keep real parameters on the real axis. This is most likely due to the failure of the simulation method, as discussed in Section D D. Additionally, due to the limited search grid of $\lambda$ used, we did not capture the entire RG flow.
Figure S4: The evolution of $\lambda_{w}=x+y\cdot g_{\varepsilon^{\prime}}$ for L=1624L=16-24 where $x$ and $y$ are fixed such that real theories reside and remain on the real axis. The points corresponding to $\lambda_{c}=0.0788\pm 0.0603$ and the evolution of values on the real axis are in blue and red respectively.
C D. Open Boundary Conditions
@@
Figure S5: The entanglement spectrum of the ground state in the (a) zero sector and (b) a charged sector of $\mathbb{Z}_{5}$ for the deformed 5-state Potts Hamiltonian at $\lambda_{c}=0.0788+0.0603i$ for system sizes L=16128L=16-128 and open boundary conditions. Here we use DMRG with $\chi=600$. The dashed lines and red numbers are the theoretical energy levels and their degeneracies, respectively [7]. We fit only the first two energy levels of the zero sector to their theoretical values.
D E. Sources of error
A large source of errors still arises from the fixed point accuracy. As discussed in Section G G, even though the precision of the fixed point was increased by an additional digit with respect to previous research [8], the repulsive nature of the point and its spiral RG flow punishes even small deviations from the true fixed point. Consequently, even for small deviations from the true fixed point, the energy levels will deviate from the expected scaling for large system sizes. We illustrate this in Figure S6, where the deviations from the theoretical scaling dimensions at $\lambda_{c}=0.0788+0.0603i$ are plotted for system sizes up to $L=64$. This repulsiveness is why the system sizes utilized in the predictions of the scaling dimensions are relatively small compared to what is computationally possible. This error is inherent to the nature of complex conformal field theories and can be resolved by repeating the procedure in Section G G for larger system sizes. However, as will be discussed next, simulation errors prevent this from being done reliably.
Figure S6: The real (a) and imaginary (b) scaling of the difference between the theoretical and estimated scaling dimensions Eq. (S.6) at $\lambda_{c}=0.0788+0.0603i$ for system sizes L=664L=6-64. The scaling generally agrees with the $1/L^{2}$ prediction up to large system sizes.
@@
Working within the MPS manifold leads to different sources of simulation errors. First, the MPS representation has a finite bond dimension. This introduces deviations from the true wavefunction because only finite entanglement is present. As a result, there is a perturbation from the true critical point, preventing additional precision of the fixed point [Huang2024]. Moreover, for critical systems, where the ground state’s correlation length diverges, and for excited states, which may violate the area law of entanglement, this approximation can be particularly problematic. The limited entanglement at the MPS boundary also introduces additional boundary effects. However, we confirmed that DMRG and QPA reach their bond saturation within the algorithmic tolerance for $\chi\geq 400$ and $600$, respectively, for system sizes up to $L=24$, regardless of the additional difficulties associated with simulating non-Hermitian models.
[…]
@@
For complex conformal theories with larger non-Hermitian contributions compared to their Hermitian counterpart, such as the recently discussed $O(N>2)$ nonlinear sigma models, we do not expect this algorithm to work, as the truncation errors would become too large [yang2026]. In this case the truncation has to be done based on the true entanglement structure of the generalized density matrix. Candidates for such a method include the biorthonormal-block density-matrix renormalization group algorithm or the fidelity algorithm [Zhong2025, Yamamoto(2022)].
To recapitulate, in this work we employed the standard DMRG and the Quasiparticle Ansatz to determine the ground state and excited states, respectively [9, 2]. We found these to agree with ED up to the imposed algorithmic tolerance of $10^{-6}$ for $\lambda_{c}=0.0788+0.0603i$ and $L=10$. Only when increasing the imaginary part of $\lambda$ in the deformed Potts Hamiltonian do we observe that these algorithms make errors that are multiple orders of magnitude larger than the imposed tolerance. This might be due to the smallest real ordering, used in both algorithms, no longer representing the ground state. Moreover, these algorithms do not account for the differences between left and right eigenvectors. We briefly mention that we also attempted to find excited states via DMRG by projecting out the ground state. We found these to agree with QPA for $\chi\geq 400$, but its convergence was much slower.